A certain ball always rebounds 2/3 of the distance from which it falls. If the ball is

dropped from a height of 9 feet, and later it is observed rebounding to a height of 64/81 feet, then how many times did it bounce?


1
Expert's answer
2021-10-01T14:28:30-0400

The height of the rebounding can be described as the next geometric sequence: "b(n)=9*({\frac 2 3})^{n}" , where n-number of bouncies

We have next equation:

"{\frac {64} {81}}=9*({\frac 2 3})^{x}"


"({\frac 2 3})^{x} = {\frac {64} {729}}"


"x=6"

The ball did bounce 6 times